OpenStudy (anonymous):

How do you figure out a specific term in an arithmetic sequence? For instance, the 46th term in the sequence -1.5, -1.3, -1.1, -0.9,...?

OpenStudy (campbell_st):

use the formula \[T_{n} = a + (n -1)\times d\] where n = the number of terms a = 1st term d = common difference ( what is being added to the term to get the next term)